HQ12.28 what volume of 0.100 M sodium chloride must be added to 7somLof0200Mlead(i) nitrate to precipitate all of the lead ions? O 15 ml. O B 30 mL D 300 ml. O E 600 ml Unanswered
Answer #1

The reaction for precipitation is:

Pb(NO3)2(aq) + 2NaCl(aq) —> PbCl2 (s) + 2NaNO3 (aq)

from given reaction:

mol of NaCl required = 2*mol of Pb(NO3)2 present

so,

M(NaCl)*V(NaCl) = 2*M(Pb(NO3)2)*V(Pb(NO3)2)

0.100 M * V(NaCl) = 2*0.200 M * 75.0 mL

V(NaCl) = 300 mL

Answer: D